Given a normal distribution with μ = 27 and σ = 25. If a random sample of size n = 72 is drawn, find P(27 ≤ x ≤ 29).

Round to three decimal places.
Answer #1

Solution :

Given that,

mean = = 27

standard deviation = = 25

= / n = 25 / 72 = 2.9463

= P[(27 - 27) /2.9463 ( - ) / (29 - 27) / 2.9463)]

= P(0 Z 0.68)

= P(Z 0.68) - P(Z 0)

= 0.7517 - 0.5

= 0.252

P(27 29) = 0.252